Tohban Report 2015-11-11
Tohban Reports | |
---|---|
Start Date: | 4 Nov 2015 |
End Date: | 11 Nov 2015 |
Tohban: | Jim McTiernan |
Tohban email: | jimm@ssl.berkeley.edu |
Next Tohban: | TBD |
List all reports |
Solar Activity
Some activity, including M flares. Less activity the last couple of days: How many GOES flares occurred?
Flares above B, C, M, X class were 11 21 4 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 0 12 4 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 4 2 0
There were RHESSI flares/GOES flares 36 / 36 over the time range 04-Nov-15 12-Nov-15
Memory Management
Decimation was active/vigorous for the full time period. Both front and rear detectors are off during SC night, except for 4 minutes on either side. The SSR is at 34% at the end of the last BGS pass this morning.
Spacecraft Status
The cold plates, 1 and 2 were at 136.7 and 135.3 K, on 2015-11-09.
Data Gaps
VC1 Gaps GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 GAP START TIME GAP END TIME GAP (SEC) 2015-11-08T09:30:00.000 -- 2015-11-08T09:35:00.000 300.00000 2015-11-08T11:10:00.000 -- 2015-11-08T11:25:00.000 900.00000
GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4 GAP START TIME GAP END TIME GAP (SEC) 2015-11-08T11:10:00.000 -- 2015-11-08T11:20:00.000 600.00000 2015-11-06T06:20:00.000 -- 2015-11-06T07:05:00.000 2700.0000
VC3 Gaps
GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30
GAP START TIME GAP END TIME GAP (SEC)
2015-11-08T02:25:00.000 -- 2015-11-08T03:55:00.000 5400.0000
Detector issues
Detector 2 is unsegmented. Detector 4 is not recording events. Detector 9 is off. Multiple fast threshold changes were made for D1 and D7 rear detectors, culminating in both D1r and D7r having rear fast thresholds of 0xFF. D7r live time has not recovered, and is currently about 5%. D1 is now at approximately 80%. On 8-nov-2015, the front fast threshold for D3 was raised to 0xF0. The was temporarily successful at raising the live time to 70%, ubt it has now dropped to 30%.
D1 and D7 live times are down to the 90-92% range, and the thin attenuator is in full time due to the low live time. The fast thresholds will be raised today o try and increase live time. Further manipulation, perhaps of the HV settings may happen today.
Other notes
Spacecraft Management
Decimation | Active/Vigorous |
HLAT Decimation | Rear decimation weight 6, no front decimation |
Night time data (fronts) | +/- 4 minutes |
Night time data (rears) | +/- 4 minutes |
Require extra passes? | No |
Requirement for moving pointer? | No |
Attenuator operation | Thin attenuator is in, see above |
Detector problems? | See notes above. |